Can someone help me on this problem?
I need to solve the followig word problem. I also need to show the equation you use for the solution.
The length of a rectangular playing field is 5 ft. less than twice its width. If the perimeter of the playing field is 230 ft., find the length and width of the field.
Thanks,
Ashley

Answer by ankor@dixie-net.com(22740)

The length of a rectangular playing field is 5 ft. less than twice its width. If the perimeter of the playing field is 230 ft., find the length and width of the field.
:
Write this sentence as an equation:
"The length is 5 ft less than twice its width."
L = 2*W - 5
:
The perimeter equation: 2L + 2W = 230
:
From the 1st statement, we know that L = (2W-5)
Substitute (2W-5) for L in the perimeter equation and solve for W:
:
2(2W-5) + 2W = 230
:
4W - 10 + 2W = 230
:
6W = 230 + 10
:
W = 240/6
:
W = 40 ft
:
I'll let you find the Length (L), check the values in the Perimeter equation